A bank quotes you an interest rate of 12% per annum with monthly compounding. What is the equivalent rate with continuous compounding?

12.22%

11.94%

11.85%

12.12%
Answer #1

EAR = (1+APR/n)^n - 1

n = number of compounds

=>

EAR = (1+12%/12)^12 - 1

= 0.12682503013

=>

e^i - 1 = 0.12682503013

=>

equivalent rate = 11.94%